OpenStudy (anonymous):

the lateral area of a right cylinder cone is 20 pi cm^2 and total area is 36pi cm^2. Find the volume of the cone.

OpenStudy (anonymous):

@sillynikki1

OpenStudy (anonymous):

@Luigi0210

OpenStudy (anonymous):

@mathmate

OpenStudy (anonymous):

@jamaica25

OpenStudy (mathmate):

 Use the fact that total area - lateral area = \(\pi r L + \pi r^2 - \pi r L = \pi r^2\) then \(\pi r^2 = 36\pi-20\pi = 16\pi\) Solve for r. Then substitute r into formula for lateral area and solve for L. Use Pythagoras theorem to find height h where \(h^2=L^2-r^2\) Finally, find volume of a cone using V = \(\pi r^2h/3\)

OpenStudy (anonymous):

thanks

OpenStudy (mathmate):

You're welcome!